# Graph sub-sampling for divide-and-conquer algorithms in large networks

### v2

@@ Abstract
As networks continue to increase in size, current methods must be capable of handling large numbers of nodes and edges in order to be practically relevant. Instead of working directly with the entire (large) network, analyzing sub-networks has become a popular approach. Due to a network’s inherent inter-connectedness, sub-sampling is not a trivial task. While this problem has gained attention in recent years, it has not received sufficient attention from the statistics community. In this work, we provide a thorough comparison of seven graph sub-sampling algorithms by applying them to divide-and-conquer algorithms for community structure and core-periphery (CP) structure. After discussing the various algorithms and sub-sampling routines, we derive theoretical results for the mis-classification rate of the divide-and-conquer algorithm for CP structure under various sub-sampling schemes. We then perform extensive experiments on both simulated and real-world data to compare the various methods. For the community detection task, we found that sampling nodes uniformly at random yields the best performance. For CP structure on the other hand, there was no single winner, but algorithms which sampled core nodes at a higher rate consistently outperformed other sampling routines, e.g., random edge sampling and random walk sampling. The varying performance of the sampling algorithms on different tasks demonstrates the importance of carefully selecting a sub-sampling routine for the specific application.
@@ 1 Introduction
Graphs, or networks [^1] We will use these terms interchangeably in this work., provide a simple and intuitive model for interconnected systems where objects or entities are represented as nodes and their relationships are denoted as edges. Graphs have been applied to many real-world applications, ranging from friendships (Girvan and Newman,, 2002) to brain connectivity (Telesford et al.,, 2011). As their popularity has grown, so too has their size: it is not uncommon for researchers to now work with thousands, millions, or even billions of nodes and edges (e.g., Backstrom et al.,, 2006; Rozemberczki et al.,, 2019). Thus, modern network analysis methods must scale to such large datasets if they are to be practically useful.
@@
One popular approach for working with large data sets in general, and large graphs in particular, is sub-sampling. The idea is simple: first the researcher divides the data into smaller sub-sets before performing analyses on the sub-sets. The analyses are then combined to give a result on the original (larger) data set. Generally, performing an analysis multiple times on smaller data sets is much faster than applying it once to the entire dataset. While sub-sampling and divide-and-conquer approaches have found great success across various statistical, machine learning and computer science tasks (Jordan,, 2012), there are some unique challenges to applying these techniques to graph-valued data. The main difficulty is that the nodes and edges which make up the graph are, by their very nature, dependent. As a graph can have a highly complex topology, it is unclear if these features will be preserved in the sub-graphs, and if not, it is unclear how that affects the analysis. Thus, it is paramount to understand the effect of the graph sub-sampling routine on the task of interest.
@@
Despite its importance, graph sampling is still an open-problem, and, in particular, has received less attention from the statistics community. Leskovec and Faloutsos, (2006) study the similarity between various measures computed on the original graph and their sub-graphs, including: clustering coefficients, degree distributions, largest eigenvectors and more. The authors find that while edge sampling performs poorly and random node sampling performs surprisingly well, random walk and forest fire samplers yield sub-networks most similar to the original network. Recently, Rozemberczki et al., (2020) develop a package to easily implement various sub-sampling routines. The new package is then applied to a variety of tasks, from computing average degrees to node classification. To our knowledge, their are not any works studying the effects of graph sub-sampling from a statistical angle, nor specifically studying their effect on divide-and-conquer algorithms.
@@
In this work, we seek to close this gap by comparing graph sub-sampling methods on divide-and-conquer algorithms. We focus on two of the most important meso-scale structures: community structure (Newman,, 2006), where nodes cluster into highly connected groups, and core-periphery structure (Yanchenko and Sengupta,, 2023), where nodes are either in the densely-connected core or sparsely-connected periphery. Our main contribution is a thorough comparison, both theoretically and empirically, of seven sub-sampling routines on the performance of divide-and-conquer algorithms for identifying these structures. After introducing the methods, we recap the theoretical results from Mukherjee et al., (2021) before deriving novel theoretical results for the CP method of Yanchenko, (2022). In particular, we derive the general mis-classification rate of the algorithm and give specific results for several sub-sampling methods. Finally, we apply these methods to various synthetic and real-world networks. We find that random node and random node neighbor yield the best community structure identification results, while routines which sample core nodes with a high probability, e.g., edge sampling and random walk, perform the best for the CP identification task. Indeed, the sub-sampling routines that perform well for one task, generally did not perform as well on the other, underscoring the importance of carefully choosing the sub-sampling algorithm for the specific problem.
@@ 2 Methodology
We begin by discussing the divide-and-conquer algorithms for community structure and CP structure, followed by presenting seven graph sub-sampling algorithms.
@@ 2.1 Divide-and-conquer for community structure
Mukherjee et al., (2021) propose $\mathsf{PACE}$, a divide-and-conquer algorithm to identify community structure in networks. The algorithm randomly samples sub-graphs, and then applies a community detection algorithm to each sub-graph. This step repeats many times before the more challenging step of “stiching” together the results from the sub-samples. Since community labels are not identifiable, this combining step is non-trivial. To avoid this problem, the authors propose an estimate of the clustering matrix $C$ instead of directly estimating the community labels as the clustering matrix is unique. This clustering matrix can then be used to extract the community labels.
More specifically, let $A$ be the $n\times n$ adjacency matrix corresponding to graph $G$ with $m$ total edges. [^2] In this work, we only consider undirected, unweighted networks without self-loops, but the ideas can easily be generalized. Let $Z\in\{0,1\}^{n\times K}$ correspond to the true community labels where $Z_{ik}=1$ if node $i$ is in cluster $k$ for $k=1,\dots,K$ communities. We assume that $K$ is known. We can define $C=ZZ^{T}\in\{0,1\}^{n\times n}$ as the clustering matrix where $C_{ij}=1$ if nodes $i$ and $j$ are in the same community and 0 otherwise. Notice that $C$ does not depend on the exact community labels (which are equivalent up to permutation) but only on the pair-wise relationship between nodes.
$\mathsf{PACE}$ begins by randomly sampling a sub-graph $S_{b}$ of graph $G$ with $qn$ nodes where $q\in(0,1)$ is the proportion of nodes in the sub-sample. With sub-sample in hand, we can apply any community detection algorithm to this sub-sample, e.g., spectral (Ng et al.,, 2002), modularity maximization (Girvan and Newman,, 2002), semi-definite programming (Cai and Li,, 2015), etc. The resulting membership vectors of these clustering algorithms will allow us to construct $\hat{C}^{(b)}\in\{0,1\}^{n\times n}$ where $\hat{C}^{(b)}_{ij}=1$ if both nodes $i$ and $j$ appear in sub-graph $S_{b}$ and were assigned to the same community and 0 otherwise. We repeat this process for $B$ sub-graphs, yielding $\hat{C}^{(1)},\dots,\hat{C}^{(B)}$.
The key step in this algorithm is combining the results from the $B$ sub-graphs. Let $N_{ij}$ be the number of times that both nodes $i$ and $j$ were chosen in the same sub-sample. Then the combined estimator $\hat{C}\in[0,1]^{n\times n}$ of the clustering matrix is defined by
@@
where $1\leq\beta\leq B$ is a tuning parameter. The combined estimator $\hat{C}$ is the proportion of sub-samples where nodes $i$ and $j$ were both sampled and assigned to the same community. We compute this quantity, however, only for the node pairs which were sampled a sufficient number of times ($>\beta$ times), otherwise the estimate is set to 0. Lastly, we can apply a clustering algorithm, e.g., $k$-means, to $\hat{C}$ in order to obtain the final community membership vector. The full algorithm is reported in Algorithm 1.
Result: Community membership vector $\hat{\boldsymbol{c}}$
Input: $n\times n$ adjacency matrix $A$, number of communities $K$, proportion of nodes to sub-sample $q$, number of sub-samples drawn $B$, sub-sampling scheme $\mathcal{S}$, clustering algorithm $\mathcal{A}$, tuning parameter $\beta$
@@
       Randomly sample sub-network $S^{(b)}$ using $\mathcal{S}$  Apply $\mathcal{A}$ to $S^{(b)}$  Obtain clustering matrix $\hat{C}^{(b)}\in\{0,1\}^{qn\times qn}$
@@
Compute $N$ where $N_{ij}$ is the number of times nodes $i$ and $j$ were sampled together in a subgraph.  Compute $\hat{C}$ where $\hat{C}_{ij}=\mathbb{I}(N_{ij}>\beta)\frac{\sum_{b=1}^{B}\hat{C}_{ij}^{(b)}}{N_{ij}}$  Perform $k$-means clustering on $\hat{C}$ to get final community membership vector $\hat{\boldsymbol{c}}\in\{1,\dots,K\}^{n}$
@@
The key advantage of this algorithm is that it estimates $C$, the clustering matrix, instead of directly estimating the community memberships $Z$, as the clustering matrix is independent of label permutations. Moreover, the estimate of $\hat{C}_{ij}$ is intuitive, namely the number of times nodes $i$ and $j$ were assigned to the same community divided by the number of times they were sampled together, given they were sampled together a sufficient number of times. Checking if $N_{ij}>\beta$ acts as a smoothing step by discarding the estimates for node pairs which only appeared in a few sub-graphs. Of course, $\beta$ is a tuning parameter that must be chosen by the user. As suggested by the authors, we set $\beta$ to be the 40th percentile of $N_{ij}$.
The main focus of this paper is the sub-sampling step of the divide-and-conquer algorithm, a topic that the authors of $\mathsf{PACE}$ give some attention too (see Section 2.3 of Mukherjee et al., (2021)). In particular, they suggest randomly sampling nodes, finding $h$-hop neighbors, onion neighborhoods or sub-graphs with roots at high degree nodes as possible sub-sampling schemes, but for nearly all experiments, they simply use random node sampling. Finally, note that Mukherjee et al., (2021) also propose another divide-and-conquer algorithm, $\mathsf{GALE}$, but for simplicity we will only consider $\mathsf{PACE}$. Indeed, in general the authors showed the superior performance of $\mathsf{PACE}$ compared to $\mathsf{GALE}$, as well as $\mathsf{GALE}$ performing best with random node sampling.
@@ 2.2.1 Core-periphery metric
Before presenting the divide-and-conquer algorithm, we first discuss the metric from Borgatti and Everett, (2000) (BE) used to quantify the CP structure of a network. The authors find the correlation between the observed network and a network with “ideal” CP structure. Let $A$ again be the adjacency matrix for the observed network and $\boldsymbol{c}\in\{0,1\}^{n}$ be the CP labels where $c_{i}=1$ if node is in the core, and 0 otherwise. Then the BE metric is
$\rho(A,\boldsymbol{c})=\mathsf{Cor}(A,\Delta_{\boldsymbol{c}})$ (2)
where $\mathsf{Cor}(A,B)$ is the Pearson correlation of the vectorized upper-triangles of matrices $A$ and $B$, and $\Delta_{\boldsymbol{c}}$ represents the ideal CP structure with $(\Delta_{\boldsymbol{c}})_{ij}=c_{i}+c_{j}-c_{i}c_{j}$. In words, $(\Delta_{\boldsymbol{c}})_{ij}=1$ if either node $i$ or $j$ is in the core, and 0 otherwise. The metric in (2) can be optimized using the greedy algorithm proposed in Yanchenko, (2022) to approximate
$\tilde{\boldsymbol{c}}=\arg\max_{\boldsymbol{c}\in\{0,1\}^{n}}\mathsf{Cor}(A,\Delta_{\boldsymbol{c}}).$ (3)

### v3

@@ Abstract
As networks continue to increase in size, current methods must be capable of handling large numbers of nodes and edges in order to be practically relevant. Instead of working directly with the entire (large) network, analyzing sub-networks has become a popular approach. Due to a network’s inherent inter-connectedness, however, sub-sampling is not a trivial task. While this problem has gained popularity in recent years, it has not received sufficient attention from the statistics community. In this work, we provide a thorough comparison of seven graph sub-sampling algorithms by applying them to divide-and-conquer algorithms for community structure and core-periphery (CP) structure. After discussing the various algorithms and sub-sampling routines, we derive theoretical results for the mis-classification rate of the divide-and-conquer algorithm for CP structure under various sub-sampling schemes. We then perform extensive experiments on both simulated and real-world data to compare the various methods. For the community detection task, we found that sampling nodes uniformly at random yields the best performance, but that sometimes the base algorithm applied to the entire network yields better results both in terms of identification and computational time. For CP structure on the other hand, there was no single winner, but algorithms which sampled core nodes at a higher rate consistently outperformed other sampling routines, e.g., random edge sampling and random walk sampling. Unlike community detection, the CP divide-and-conquer algorithm tends to yield better identification results while also being faster than the base algorithm. The varying performance of the sampling algorithms on different tasks demonstrates the importance of carefully selecting a sub-sampling routine for the specific application.
Keywords: Exploration sampling; Community detection; Core-periphery; Mis-classification rate; Network analysis
@@ 1 Introduction
Graphs, or networks [^1] We will use these terms interchangeably in this work., provide a simple and intuitive model for interconnected systems where objects or entities are represented as nodes and their relationships are denoted as edges. Networks have been applied to many real-world applications, ranging from friendships (Girvan and Newman,, 2002) to brain connectivity (Telesford et al.,, 2011). As their popularity has grown, so too has their size: it is not uncommon for researchers to now work with thousands, millions or even billions of nodes and edges (e.g., Backstrom et al.,, 2006; Rozemberczki et al.,, 2021). Community detection (Mukherjee et al.,, 2021; Fang et al.,, 2024; Cardoso et al.,, 2024; Chakrabarty et al.,, 2025), hypothesis testing (Ghoshdastidar and Von Luxburg,, 2018; Chakraborty et al.,, 2024) and link prediction (Zhang and Chen,, 2018; Kumar et al.,, 2020) are just a few examples where researchers must handle large graphs. Thus, modern network analysis methods must scale to such datasets if they are to be practically useful.
@@
One popular approach for working with large data sets in general, and large graphs in particular, is sub-sampling. The idea is simple: first the researcher divides the data into smaller sub-sets before performing analyses on the sub-sets. The analyses are then combined to give a result on the original (larger) data set. Generally, performing an analysis multiple times on smaller data sets is much faster than applying it once to the entire dataset. While sub-sampling and divide-and-conquer approaches have found great success across various statistical, machine learning and computer science tasks (Jordan,, 2012), there are some unique challenges to applying these techniques to graph-valued data. The main difficulty is that the nodes and edges which make up the graph are, by their very nature, dependent. As a graph can have a highly complex topology, it is unclear if these features will be preserved in the sub-graphs, and if not, it is unclear how this affects the analysis. Thus, it is paramount to understand the effect of the graph sub-sampling routine on the task of interest.
@@
Despite its importance, graph sampling is still an open-problem, and, in particular, has received less attention from the statistics community. Leskovec and Faloutsos, (2006) study the similarity between various measures computed on the original graph and their sub-graphs, including: clustering coefficients, degree distributions, largest eigenvectors and more. The authors find that while edge sampling performs poorly and random node sampling performs surprisingly well, random walk and forest fire samplers yield sub-networks most similar to the original network. Instead of comparing different sub-sampling strategies, other works focus their study on a single paradigm, such as random walks (e.g., Ribeiro et al.,, 2012; Li et al.,, 2015; Chiericetti et al.,, 2016)
Several recent works have proposed novel graph sampling algorithms for various tasks. Wang, (2023) develop an adaptive edge sub-sampling algorithm and apply it to find degree, clustering coefficient and K-core distributions. Similarly, Chen and Su, (2024) propose an edge-reinforced random walk sub-sampling strategy and show that sub-graphs extracted using the algorithm are similar to the original network, while Wang et al., (2024) propose a motif-based sub-sampler for supports estimation.
Recently, Rozemberczki et al., (2020) develop a package to easily implement various sub-sampling routines. The new package is then applied to a variety of tasks, from computing average degrees to node classification. Rožanec and Rožanec, (2024) provide a similar package implementing various sub-sampling algorithms in the Go programming language. To our knowledge, there are not any works studying the effects of graph sub-sampling from a statistical angle, nor specifically studying their effect on divide-and-conquer algorithms.
@@
In this work, we seek to close this gap by comparing graph sub-sampling methods on divide-and-conquer algorithms. We focus on two of the most important meso-scale structures: community structure (Newman,, 2006), where nodes cluster into highly connected groups, and core-periphery structure (Yanchenko and Sengupta,, 2023), where nodes are either in the densely-connected core or sparsely-connected periphery. Our main contribution is a thorough comparison, both theoretically and empirically, of seven sub-sampling routines on the performance of divide-and-conquer algorithms for identifying these structures. After introducing the methods, we recap the theoretical results from Mukherjee et al., (2021) before deriving novel theoretical results for the CP method of Yanchenko, (2022). In particular, we derive the general mis-classification rate of the algorithm and give specific results for several sub-sampling methods. Finally, we apply these methods to various synthetic and real-world networks. We find that random node and random node neighbor yield the best community structure identification results, while routines which sample core nodes with a high probability, e.g., edge sampling and random walk, perform the best for the CP identification task. Indeed, the sub-sampling routines that perform well for one task, generally did not perform as well on the other, underscoring the importance of carefully choosing the sub-sampling algorithm for the specific problem. Moreover, the performance of the base algorithm relative to the divide-and-conquer approach also differed depending on the task. Notably, the CP divide-and-conquer algorithm outperformed the base algorithm both in terms of detection ability and computation time.
@@ 2 Methodology
We begin by discussing the divide-and-conquer algorithms for community and CP structure, followed by presenting seven graph sub-sampling algorithms.
2.1 Notation
Let $G=(V,E)$ be a graph with $|V|=n$ nodes and $|E|=m$ edges. We also define ${\bf A}$ as the $n\times n$ adjacency matrix corresponding to graph $G$ such that $A_{ij}=1$ if nodes $i$ and $j$ are connected by and edge, and 0 otherwise. [^2] In this work, we only consider undirected, unweighted networks without self-loops, but the ideas can easily be generalized. Let $K$ be the known (integer) number of communities, and let ${\bf Z}\in\{0,1\}^{n\times K}$ correspond to the true community labels where $Z_{ik}=1$ if node $i$ is in community $k$ for $k=1,\dots,K$. We define ${\bf C}={\bf Z}{\bf Z}^{T}\in\{0,1\}^{n\times n}$ as the clustering matrix where $C_{ij}=1$ if nodes $i$ and $j$ are in the same community, and 0 otherwise. Notice that ${\bf C}$ does not depend on the exact community labels (which are equivalent up to permutation) but only on the pair-wise relationship between nodes.
@@ 2.2 Divide-and-conquer for community structure
Mukherjee et al., (2021) propose $\mathsf{PACE}$, a divide-and-conquer algorithm to identify community structure in networks. The algorithm randomly samples sub-graphs, and then applies a community detection algorithm to each sub-graph. This step repeats many times before the more challenging step of “stitching” together the results from the sub-samples. Since community labels are not uniquely identifiable, this combining step is non-trivial. To avoid this problem, the authors estimate the clustering matrix ${\bf C}$ instead of directly estimating the community labels as the clustering matrix is unique. This clustering matrix can then be used to extract the community labels.
$\mathsf{PACE}$ begins by randomly sampling a sub-graph $S^{(b)}$ of graph $G$ with $qn$ nodes where $q\in(0,1)$ is the proportion of nodes in the sub-sample. With a sub-sample in hand, we can apply any community detection algorithm to this sub-sample, e.g., spectral (Ng et al.,, 2002), modularity maximization (Girvan and Newman,, 2002), semi-definite programming (Cai and Li,, 2015), etc. The resulting membership vectors of these clustering algorithms allow us to construct $\hat{\bf C}^{(b)}\in\{0,1\}^{n\times n}$ where $\hat{C}^{(b)}_{ij}=1$ if both nodes $i$ and $j$ appear in sub-graph $S^{(b)}$ and were assigned to the same community, and 0 otherwise. We repeat this process for $B$ sub-graphs, yielding $\hat{\bf C}^{(1)},\dots,\hat{\bf C}^{(B)}$.
The key step in this algorithm is combining the results from the $B$ sub-graphs. Let $N_{ij}$ be the number of times that both nodes $i$ and $j$ were chosen in the same sub-sample. Then the combined estimator $\hat{\bf C}\in[0,1]^{n\times n}$ of the clustering matrix is defined by
@@
where $\mathbb{I}(\cdot)$ is the indicator function, and $1\leq\beta\leq B$ is a tuning parameter. The combined estimator $\hat{\bf C}$ is the proportion of sub-samples where nodes $i$ and $j$ were both sampled and assigned to the same community. We only compute this quantity, however, for node pairs which were sampled a sufficient number of times ($>\beta$ times), otherwise the estimate is set to 0. Lastly, we can apply a clustering algorithm, e.g., $k$-means, to $\hat{\bf C}$ in order to obtain the final community membership vector. The full algorithm is reported in Algorithm 1.
Result: Community membership vector $\hat{\boldsymbol{c}}\in\{1,\dots,K\}^{n}$
Input: graph $G$, number of communities $K\in\mathbb{Z}^{+}$, proportion of nodes to sub-sample $q\in(0,1)$, number of sub-samples $B\in\mathbb{Z}^{+}$, sub-sampling scheme $\mathcal{S}$, clustering algorithm $\mathcal{A}$, tuning parameter $\beta\in\{1,2,\dots,B\}$
@@
       Randomly sample $G$ using $\mathcal{S}$, resulting in sub-network $S^{(b)}$  Apply $\mathcal{A}$ to $S^{(b)}$  Obtain clustering matrix $\hat{\bf C}^{(b)}\in\{0,1\}^{qn\times qn}$
@@
Compute ${\bf N}$ where $N_{ij}$ is the number of times nodes $i$ and $j$ were sampled together in a subgraph.  Compute $\hat{\bf C}$ where $\hat{C}_{ij}=\mathbb{I}(N_{ij}>\beta)\frac{\sum_{b=1}^{B}\hat{C}_{ij}^{(b)}}{N_{ij}}$  Perform $k$-means clustering on $\hat{\bf C}$ with $K$ clusters to obtain $\hat{\boldsymbol{c}}\in\{1,\dots,K\}^{n}$
@@
The key advantage of this algorithm is that it estimates ${\bf C}$, the clustering matrix, instead of directly estimating the community memberships ${\bf Z}$, as the clustering matrix is independent of label permutations. Moreover, the estimate of $\hat{C}_{ij}$ is intuitive, namely the number of times nodes $i$ and $j$ were assigned to the same community divided by the number of times they were sampled together, given they were sampled together a sufficient number of times. Checking if $N_{ij}>\beta$ acts as a smoothing step by discarding the estimates for node pairs which only appeared in a few sub-graphs. Of course, $\beta$ is a tuning parameter that must be chosen by the user. As suggested by the authors of $\mathsf{PACE}$, we set $\beta$ to be the 40th percentile of $N_{ij}$. Additionally, note that the final $k$-means clustering step could fail by providing zero- or one-node clusters, which is a limitation of this method.
The main focus of this paper is the sub-sampling step of the divide-and-conquer algorithm, a topic that the authors of $\mathsf{PACE}$ give some attention too (see Section 2.3 of Mukherjee et al., (2021)). In particular, as possible sub-sampling schemes, they suggest: randomly sampling nodes; randomly sampling a node and including all of its neighbors within distance $h$ for some integer $h$ ($h$-hop neighborhoods); $1$-hop neighborhoods but removing the root node (ego neighborhoods); randomly sampling a node, including all of its neighbors, including all of the neighbors of the neighbors, etc. (onion neighborhoods); or sampling sub-graphs with roots at high degree nodes. For most experiments, the authors employ random node sampling.
@@ 2.3.1 Core-periphery metric
Before presenting the divide-and-conquer algorithm, we first discuss the metric from Borgatti and Everett, (2000) (BE) used to quantify the CP structure of a network. The authors define the core as a set of nodes which are highly connected with each other, as well as with the peripheral nodes. In particular, they find the correlation between the observed network and a network with “ideal” CP structure. Let ${\bf A}$ again be the adjacency matrix for the observed network and $\boldsymbol{c}\in\{0,1\}^{n}$ be some CP labels where $c_{i}=1$ if node is assigned to the core, and 0 otherwise. Then the BE metric is
$\rho({\bf A},\boldsymbol{c})=\mathsf{Cor}({\bf A},\Delta_{\boldsymbol{c}})$ (2)
where $\mathsf{Cor}({\bf A},{\bf B})$ is the Pearson correlation of the vectorized upper triangular parts of matrices ${\bf A}$ and ${\bf B}$, and $\Delta_{\boldsymbol{c}}$ represents the ideal CP structure with $(\Delta_{\boldsymbol{c}})_{ij}=c_{i}+c_{j}-c_{i}c_{j}$. In words, $(\Delta_{\boldsymbol{c}})_{ij}=1$ if either node $i$ or $j$ is in the core, and 0 otherwise. The metric in (2) can be optimized using the greedy algorithm proposed in Yanchenko, (2022) to approximate
$\hat{\boldsymbol{c}}=\arg\max_{\boldsymbol{c}\in\{0,1\}^{n}}\mathsf{Cor}({\bf A},\Delta_{\boldsymbol{c}}).$ (3)